1) determine the index of hydrogen deficiency (HD) 2) draw the possible structures based on the IHD 3) for each structure, inThis is an organic chemistry problem and i need them numbered out and written properly and correctly. any written mistake will be thumbs down. USE THe structure as reference and Do ALL PROBLEMS ASAP
Answer #1

C₂ H3U : 10) IHD = IHO (20+2) -H-X - C = # of c. (x=() = # of ce - IHD = (2x2+2) - 3-3 THD and = 60 =0 suggests that there is3 No. of cufferent types of hydrogens can be counted simply by lij substituting all the hydrogens with any other group and seHa HC-843 Let us This is one type of H. (+) Hi relationship between Ha 4 Hb. EC-(uz (1) & (11) are enantiomers so Has Hy will